OpenStudy (anonymous):

Solve the equation. If it is an identity (all real numbers) or a contradiction (no solution), so state. 3x + 5 + 6x = 9x + 8

OpenStudy (anonymous):

3x + 5 + 6x = 9x + 8 9x + 5 = 9x + 8 9x - 9x = 8 -5 0 = 3 this is a contradiction, so no solution